OpenStudy (anonymous):

The population of a town grows at a rate proportional to the population present at time t. The initial population of 50 increases by 15% in 10 years. What will be the population in 30 years? how fast is the population growing at t=30?

OpenStudy (anonymous):

\[\frac{dP}{dt}=kA\]

OpenStudy (anonymous):

kP i mean

OpenStudy (anonymous):

\[P'-kP=0\]

OpenStudy (anonymous):

\[e^{\int{}{}p(x)dx}=e^{kt}\]

OpenStudy (anonymous):

+c of course

OpenStudy (anonymous):

actually not yet

OpenStudy (anonymous):

\[\frac{d}{dx}(e^{-kt}y)=0\]

OpenStudy (anonymous):

\[e^{-kt}y=c\]

OpenStudy (anonymous):

\[e^{-kt}P=c\]

OpenStudy (anonymous):

P(0)=500

OpenStudy (anonymous):

\[e^{-k(0)}*500=c\]= \[500=c\]

OpenStudy (anonymous):

P(10)=500+(.15*500)=575 \[e^{-10k}*500=575\]

OpenStudy (anonymous):

\[e^{-10k}=\frac{575}{500}\]

OpenStudy (anonymous):

\[-10k=ln(\frac{575}{500})\]

OpenStudy (anonymous):

\[-10k=ln(575)-ln(500)\] \[k=\frac{ln(575)-ln(500)}{-10}\]

OpenStudy (asnaseer):

I would have just done this:\[\begin{align} \frac{dP}{dt}&=kP\\ \therefore\int\frac{dP}{P}&=\int k dt\\ \therefore\log(P)&=kt+c\\ \therefore P&=e^{kt+c}=e^c\times e^{kt}=Ae^{kt}\\ \end{align}\]when t=0, P=50, therefore:\[50=Ae^0=A\]\[P=50e^{kt}\]

OpenStudy (asnaseer):

then, in 10 years the population increases by 15%, which means it becomes 50*1.15 = 57.5

OpenStudy (asnaseer):

therefore:\[57.5=50e^{10k}\]\[e^{10k}=\frac{57.5}{50}\]\[10k=\log(\frac{57.5}{50})\]\[k=\frac{1}{10}\log(\frac{57.5}{50})\]

OpenStudy (anonymous):

yeah it's 575 i mean

OpenStudy (asnaseer):

so you can now work out the P and dP/dt when t=30

OpenStudy (anonymous):

yeah how'd you get 10k though i'm wondering

OpenStudy (anonymous):

because i did it through linear and it came out negative

OpenStudy (asnaseer):

ok, you agree we got to this point earlier where k was some unknown constant?\[P=50e^{kt}\]

OpenStudy (anonymous):

yes

OpenStudy (asnaseer):

then you are told that when t=10, P=50 increased by 15%=57.5 agreed?

OpenStudy (anonymous):

yes your way but if you do it with the linear equation y'-x(t)y=g(t)

OpenStudy (anonymous):

y'+p(x)y=g(x)

OpenStudy (asnaseer):

ok, using your way you ended up with:\[e^{-kt}P=c\]if you multiply both sides by \(e^{kt}\) you get:\[P=ce^{kt}\]which is the same form as mine.

OpenStudy (anonymous):

if you leave it as so and solve

OpenStudy (anonymous):

\[e^{-10k}(575)=500\]

OpenStudy (asnaseer):

you mean leave it in your form?

OpenStudy (anonymous):

yes idk if it's just the answers are the sme

OpenStudy (anonymous):

\[e^{-10k}=\frac{500}{575}\] \[-10k=ln(\frac{500}{575}\]

OpenStudy (anonymous):

cause when you get farther you're taking the diference of two logs correct? does the -10 correct that

OpenStudy (asnaseer):

it should lead to the same answers. from your equation we get:\[k=-\frac{\log(\frac{500}{575})}{10}=-\frac{\log(500)-\log(575)}{10}=\frac{\log(575)-\log(500)}{10}\]

OpenStudy (asnaseer):

and this is identical to my result above of:\[k=\frac{1}{10}\log(\frac{57.5}{50})\]